The yellow light given off by a sodium vapour lamp used for public lighting has a wavelength of 589 nm. What is the frequency and energy of this radiation?


1
Expert's answer
2021-06-16T05:32:14-0400

Velocity = frequency × wavelength

F= c/α

= 3.0 ×108 / 589×10-9

= 5.09 × 1014 Hz